OpenStudy (starryknight):

find the unit price(dollars per ounce)Dont understand: A 20 ounce can of pineapple juice for $0.98

OpenStudy (mathmale):

You have 2 separate problems here. Please do one at a time. Better yet, post only one at a time.

OpenStudy (anglina4196):

Wait do you have to find the unit price or the dollars per ounce

563blackghost (563blackghost):

Unit price is basically like a simplified fraction.....We are given `$.98` for `20-ounces` of pineapple juice...now we want to know how much it costs for 1 ounce of pineapple juice....in order to do this we would need to divide the total price for 20-ounces of juice by 20 ounces to get the price for 1 ounce.... \(\huge{\frac{total~price}{ounces}\rightarrow\frac{.98}{20}}\) So divide....

OpenStudy (anglina4196):

divide .98 by 20

OpenStudy (mathmale):

It is essential that you include the units of measurement here. Not .98 /20, but $0.98 / 20 ounces = \[\frac{ ?~dollars }{ ounce }\]